OpenStudy (anonymous):

Solve for x. Log(x+2)+log(x-4)=log55

OpenStudy (therealmeeeee):

ok log 2 + log x = log 3 or, log x = log 3 - log 2 or, log x = log(3/2) so, x = 3/2 Eq 2 : log 2 + log x = 3 or, log(2 * x) = 3 or, 2*x = 10 ^ 3 or, x = (10 ^ 3 )/ 2 so , x = 500

OpenStudy (therealmeeeee):

ok now do this one help u

OpenStudy (anonymous):

No are you reading th question right

OpenStudy (therealmeeeee):

yes i am

OpenStudy (therealmeeeee):

rewrite it then but better

OpenStudy (anonymous):

That is a clear as it gets. The numbers your using are not right

OpenStudy (shamil98):

log(x+2)+ log(x-4) = log 55 Using the log rule.. \[\large \log_b(mn) = \log_b M + \log_b N\] \[\large \log (x+2)(x-4) = \log 55\]

OpenStudy (anonymous):

I know all the rules but that is not helping in this case.

OpenStudy (anonymous):

@Bearcaesar the reason @TheRealMeeeee 's solutions are wrong is that because he is only copy and pasting it from google. lol

OpenStudy (shamil98):

I'm getting to that.. you can eliminate it like this: \[\large \frac{ \log (x+2)(x-4) }{ \log(55) } = \frac{ \log 55 }{ \log 55 }\]

OpenStudy (shamil98):

The resulting equation is.. \[\large \frac{ (x+2)(x-4) }{ 55 } = 1\] Solve it algebraically from here.

OpenStudy (anonymous):

Thanks GirlByte. I was like what is going on.

OpenStudy (anonymous):

Thanks so much shamil98

OpenStudy (shamil98):

\[\large (x+2)(x-4) = 55\] \[\large x^2 -2x - 8 = 55\] and so on, and no problem :)

OpenStudy (anonymous):

Wait but that does not give me the answer because I did it that way

OpenStudy (shamil98):

It does.. \[\huge x^2 - 2x - 63 = 0\]

OpenStudy (anonymous):

Believe me I entered that answer at least eight times. I am supposed to have x equals something.

OpenStudy (shamil98):

Your supposed to solve for x after wards .... \[\huge (x+7)(x-9) = 0\] \[\huge x = 9 , x = -7\] x = -7 back into the original equation this solution does not work so your only solution is \[\Huge x=9\]

OpenStudy (anonymous):

Ohhhh makes sense

OpenStudy (shamil98):

mhm.